OpenStudy (anonymous):

What is the equation for the vertical asymptote of the function shown below? f(x)=3x^4-3/2x-5 A. x = 3/5 B. y = x3 - 2 C. x = 3/2 D. x = 5/2

OpenStudy (amistre64):

vas are where the final factored form creates a zero in the denominator

OpenStudy (anonymous):

so i need to factor \[f(x)=3x^4-3/2x-5\]

OpenStudy (amistre64):

3x^4-3 3(x^4-1) ------- = -------- ... i dont see any factors canceling soo 2x-5 2x-5 2x-5 = 0 whent = x 5/2 right?

OpenStudy (amistre64):

factoring helps to weed out the holes from the vas

OpenStudy (amistre64):

holes are where the graph jumps of a spot; vas are where they slide up real close to it and shoot off into infinity

OpenStudy (amistre64):

..jumps over a spot ....

OpenStudy (anonymous):

so i should plug in thte answers they gave me?

OpenStudy (amistre64):

B is obviously not a VA if we stick to standard conventions of placements for the x and y axises

OpenStudy (anonymous):

or is the answer just 5/2?

OpenStudy (amistre64):

Look for numbers that zero out the bottom; those are possible candidates for VAs

OpenStudy (amistre64):

if the top and bottom factor and cancel out like terms; the ones that are left are VAs .... this one has no like terms to cancel, so anyting that zeros out the bottom is a VA

OpenStudy (amistre64):

3/2 and 3/5 do not produce a zero for the bottom 5/2 does; so the VA is x=5/2

OpenStudy (anonymous):

oooh i see thanks!

OpenStudy (amistre64):

youre welcome